# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

### v1

@@ 8 Summary
In this work, we propose OpeNLGauge – a versatile method for evaluating NLG that uses only open-weight models and provides fine-grained explainability. The method provides a much better explanation quality than previous methods and achieves competitive correlations with human judgments.
@@
OpeNLGauge is a method for evaluating a variety of NLG tasks. While this paper presents the evaluation on several NLG tasks and the method achieves good performance on unseen aspects, domains and tasks, the actual performance on new NLG tasks is unknown. In particular, the metric has not been tested in a multilingual setting. Moreover, previous research has shown that some LLM-based metrics have a bias towards texts generated by LLMs Liu et al. (2023).
@@ Acknowledgments
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@ Appendix A Aspects and datasets in the training data
This section provides a detailed overview of source datasets and evaluation aspects used to generate our training dataset.
@@ A.1 Summarization
As source data for summarization, we utilize five datasets that include three distinct tasks: news article summarization, forum post summarization and dialogue summarization. Our evaluation dataset includes six commonly used aspects: four related to meaning of the evaluated text, one to its form and one to both. Some of these aspects overlap significantly in their definitions. This is intentional, as our goal is to make the evaluator model robust to small variations in aspect definitions, and enable generalization to new aspects. Table 6 provides an overview of the evaluated systems.
@@ A.1.1 Datasets
  • CNN/DailyMail (Hermann et al., 2015) is a popular summarization dataset that consists of news articles from CNN and DailyMail, paired with corresponding bullet-point summaries. Originally developed for question answering, it was later adapted for summarization (Nallapati et al., 2016). For this dataset, we use pre-generated outputs from Stiennon et al. (2020), which include results from 11 different systems, including human references and an extractive baseline. Since our meta-evaluation datasets contain inputs from CNN/DailyMail, we ensure there is no overlap in the source texts between these datasets when sampling the inputs.
  • Newsroom (Grusky et al., 2018) is a large-scale dataset of news articles and their summaries, collected from diverse sources, domains, authors and time range. In addition to newly generated LLM outputs, we use outputs of the systems evaluated in the original paper, which include three summarization systems and two extractive baselines.
  • SAMSum dataset (Gliwa et al., 2019) addresses the dialogue summarization task, containing dialogues with short summaries created by linguists. In addition to outputs from several different LLMs, we use pre-generated outputs of six systems from (Gao and Wan, 2022).
  • TL;DR (Völske et al., 2017) is a collection of Reddit posts with user-created summaries. Unlike many popular datasets that focus on news articles, TL;DR contains informal and less structured texts spanning diverse topics. Similarly to CNN/DailyMail, we use outputs from Stiennon et al. (2020).
  • XSum (Narayan et al., 2018) is a dataset for extreme summarization task, designed for abstractive approaches. Unlike datasets such as Newsroom or CNN/DailyMail, which favor extractive summarization, XSum contains BBC articles paired with concise, single-sentence summaries. We utilize pre-generated outputs from five systems and baselines evaluated in the original paper, along with newly generated output from LLMs.
@@ A.1.2 Aspects
  • Accuracy is largely synonymous with consistency. It evaluates whether the factual information from the source is accurately represented in the summary. Following Stiennon et al. (2020), we define accuracy as: Extent to which the factual information in the summary accurately matches the source text. An accurate summary should not contain information that is not present in the source text, should not contradict the source text, and generally should not be misleading.
@@
  • Coverage evaluates how much of the important information from the source text is covered by the summary. In this sense, it is closely related to relevance, however, it does not include non-redundancy as a criterion. We use a definition adapted from Stiennon et al. (2020): Extent to which the summary covers the important information in the source text. A summary has good coverage if it mentions the main information from the source text that is important to understand the events described in the text. A summary has poor coverage if someone reading only the summary would be missing several important pieces of information about the event in the source text.
  • Coherence refers to the structural quality of the summary and involves attributes such as cohesion, consistency and relevance (Reinhart, 1980). It is determined by both semantic and formal structure of the text. We define coherence as: Extent to which the summary is well-structured and organized, presenting information in a logical order that flows naturally from sentence to sentence. Coherent summary forms a unified body of information and makes it easy to understand the main ideas.
@@
Table 5: Input data formats for data-to-text tasks.
@@ A.2 Data-to-text
For data-to-text category, we collected inputs from four datasets that represent four distinct tasks: table-to-text, RDF-to-text, attribute-value list to text and logical NLG. Since it is important for an NLG evaluation method to reliably evaluate outputs with respect to structured data in different formats, we include a number of different input formats in our training data, including JSON, CSV and linearized tables with markup. Table 5 provides an overview of the input formats. The list of evaluated systems is shown in Table 7.
@@ A.2.1 Datasets
  • E2E NLG dataset (Dušek et al., 2020; Novikova et al., 2017b) was chosen as a representative of simple data-to-text tasks. In this dataset, models are tasked with generating descriptions of restaurant venues based on attribute-value-based meaning representations (MRs). Target descriptions were crowdsourced using textual and pictorial representations of MRs as stimuli. To ensure diversity, we selected pre-generated outputs from five systems described by (Dušek et al., 2020), considering their model architectures and evaluation results across various metrics and aspects. These were extended by new outputs from several LLMs. The inputs and pre-generated outputs were sampled from the test set. Three different input formats are used in the training data (see Table 5).
  • WebNLG 2020 (Ferreira et al., 2020) is an RDF-to-text dataset designed for generating natural language text from RDF triples collected from DBPedia knowledge base (Mendes et al., 2012). Each input contains between one and seven triples, where each triple represents a binary relation in the form (subject, property, object). Similarly to E2E NLG, we selected a diverse set of output systems based on model architectures and evaluation results from the WebNLG+ 2020 Challenge, and used additional LLMs to generate new outputs.
  • ToTTo (Parikh et al., 2020) is an open-domain table-to-text generation dataset. It consists of Wikipedia tables with highlighted cells, and the task is to generate single–sentence descriptions of the data in these highlighted cells. Inputs are provided in two formats: full tables, which include indices to highlighted cells, and linearized tables, where only the highlighted data is presented in a linear order, while the structure is annotated with markup tags. As we observed that even medium-sized open-weight LLMs often struggle with hallucinations when using full tables, we restricted our evaluation to linearized tables.
  • LogicNLG (Chen et al., 2020) introduced the task of logical NLG, where models generate statements that can be logically entailed from the data in a table. This task involves various aggregations and comparisons, making it more difficult than simply transforming structured data to free-form text. Although the task explicitly requires generating five logical statements per input, we sampled between one to five statements from each generated output to increase the diversity of output lengths. Inputs were formatted as CSV with “|” as a separator. In addition to new LLM outputs, we used existing system output from Zhao et al. (2023b).
@@ A.2.2 Aspects
  • Faithfulness measures whether all information in the generated text is supported by the data, making it equivalent to precision. Similarly to factual consistency in summarization, we consider both factual and non–factual hallucinations as errors. For our evaluations, we define faithfulness as: Extent to which the information in the text is supported by the data.
@@
  • Fluency refers to the formal quality of the generated text, and includes grammaticality, naturalness and readability. Some definitions also include coherence (Ferreira et al., 2020), although coherence is usually treated as a separate aspect. For our evaluation, we define fluency as: Extent to which the text is grammatical, natural and easy to understand.
@@ A.3 Dialogue Response Generation
For dialogue response generation, we source the inputs from three dialogue datasets, focusing on open-domain non-task-oriented dialogue response generation. Table 8 provides an overview of the evaluated systems.
@@ A.3.1 Datasets
  • Wizard of Wikipedia (Dinan et al., 2018) consists of conversations grounded in one of 1365 topics and corresponding knowledge retrieved from Wikipedia. In these conversations, either participant may select the topic and initiate the discussion, although they have asymmetric roles. One participant takes on the role of the wizard, an expert with access to a topic-relevant knowledge, on which they can base their responses. The other participant acts as an apprentice, a curious learner that is eager to discuss the chosen topic. To create inputs of varying length, we randomly select a dialogue history length between two turns and the full conversation, and truncate the dialogue to this length. The last utterance is replaced by a system output, except when the reference is used as evaluated output. In addition to newly generated LLM responses, we include human responses from the original dataset in the outputs.
  • EmpatheticDialogues (Rashkin, 2018) is a dataset of dialogues grounded in emotional situations, designed to train and evaluate dialogue models on empathetic response generation. Each conversation is associated with an emotional label, where one of the participants describes a situation in which they experienced a given emotion. We sample up to five turns from each dialogue and replace the last utterance with a generated response. The emotion label is used as additional context for annotator LLMs to evaluate the appropriateness and empathy of the responses, but is excluded from the prompts used for system output generation.
  • DailyDialog (Li et al., 2017) includes conversations on various daily life topics, annotated with emotion labels and communicative intents. We include data from DailyDialog to represent diverse topics and scenarios in the training set. Alongside newly generated responses, we also collect pre-generated outputs from three sources (Gupta et al., 2019; Huang et al., 2020; Zhao et al., 2020) to represent older dialogue systems.
@@ A.3.2 Aspects
  • Interestingness is concerned with the informational value of the response, specifically whether it presents stimulating ideas, facts or opinions. Although it is one of the more subjective evaluation aspects, we define it vaguely and let the annotator models determine the criteria for interestingness. Therefore, in our dataset, we define interestingness as: Extent to which the response is interesting given the dialogue history.
  • Engagingness is closely related to interestingness but is sometimes treated as a distinct aspect (e.g., Mehri and Eskenazi, 2020a; See et al., 2019). While interestingness focuses on the context itself, engagingness emphasizes maintaining the user’s attention and encouraging them to continue with the conversation. For our purposes, engagingness is defined as: Extent to which the response captures and maintains the user’s interest, encouraging further interaction. Engaging responses contain opinions, preferences, thoughts or interesting facts.
@@ A.4 Story Generation
The NLG tasks discussed so far generally contain inputs that are relatively longer compared to the outputs. This pattern is especially common in tasks like summarization and dialogue generation, although certain data-to-text tasks also share this characteristic. To represent scenarios with short inputs and long outputs, we include story generation in the training data. Table 9 lists the evaluated systems.
@@ A.4.1 Datasets
As the source of inputs, we use WritingPrompts (Fan et al., 2018), a story generation dataset derived from Reddit’s WritingPrompts subreddit, where users submit prompts that can inspire other users to write stories. The dataset consists of a diverse range of topics, story lengths and writing styles. We reuse existing outputs from the OpenMEVA dataset (Guan et al., 2021) and generate additional outputs by four LLMs. To increase the diversity of generated stories in terms of their length, we generate the outputs with two different prompt versions, each requiring a different length of the story. The outputs are then randomly sampled from either the shorter or the longer set. As we observed a tendency of LLMs to generate a long list of errors for longer inputs, our evaluator models are instructed to limit the number of identified errors to a maximum of eight, and to prioritize the most severe ones if necessary.
@@ A.4.2 Aspects
While relevance and coherence are two commonly used aspects for story generation, there is no consensus on which other evaluation aspects are the most relevant. Inspired by social sciences, Chhun et al. (2022) propose four additional aspects, aimed at providing a complete and non-redundant set of criteria. Following their work, we adopt the aspects defined in the HANNA benchmark, using our own definitions for most of them:
  • Relevance measures the degree to which a story aligns with the given prompt (Chhun et al., 2022; Chiang and Lee, 2023a), title (Jhamtani and Berg-Kirkpatrick, 2020; Yao et al., 2019; Xie et al., 2023) or story beginning Wang et al. (2020b). In some cases, relevance also evaluates whether the story remains on-topic for its duration (e.g., Goldfarb-Tarrant et al., 2020). Since our inputs are prompts, we define relevance as: Extent to which the story is relevant to the writing prompt.
  • Coherence in story generation typically refers to logical consistency and narrative flow (e.g., Yao et al., 2019; Li et al., 2023b; Jhamtani and Berg-Kirkpatrick, 2020). Other works define coherence more vaguely, such as how much the story “makes sense” (Chhun et al., 2022) or how well its sentences “fit together” (Xie et al., 2023). For our purposes, coherence is defined as: Extent to which the story is logically consistent and coherent.
  • Engagement is a subjective and often vaguely defined aspect that evaluates how engaging the story is to the reader (e.g., Chhun et al., 2022; Li et al., 2023b). Due to its inherent subjectivity, we apply a simple definition and leave its interpretation to the evaluator models: Extent to which the story is engaging and interesting.
@@ A.5 Question Answering
The question answering subset of our dataset consists of two distinct tasks: narrative question answering and table question answering. The inputs consist of a question and structured data in which the answer should be grounded. The evaluated systems are listed in Table 10.

### v3

@@ 8 Summary
In this work, we present OpeNLGauge – a versatile method for evaluating NLG that uses only open-weight models and provides fine-grained explainability. The method provides a much better explanation quality than previous methods and achieves competitive correlations with human judgments.
@@
OpeNLGauge is a method for evaluating a variety of NLG tasks. While this paper presents the evaluation on several NLG tasks and the method achieves good performance on unseen aspects, domains and tasks, the actual performance on new NLG tasks is unknown. In particular, the metric has not been tested in a multilingual setting. Moreover, previous research has shown that some LLM-based metrics have a bias towards texts generated by LLMs liu-etal-2023-g.
@@ Acknowledgments
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@ Appendix A Aspects and datasets in the training data
This section provides a detailed overview of all source datasets and evaluation aspects used to generate our training data for OpeNLGaugeft.
@@ A.1 Summarization
As source data for summarization, we utilize five datasets spanning three distinct tasks: news article summarization, forum post summarization and dialogue summarization. Our evaluation dataset includes six commonly used aspects: four related to meaning of the evaluated text, one to its form and one to both. Some of these aspects overlap significantly in their definitions. This is intentional, as our goal is to make the evaluator model robust to small variations in aspect definitions, and enable generalization to new aspects. Table 6 provides an overview of all summarization systems in our dataset.
@@ A.1.1 Datasets
  • CNN/DailyMail (hermann2015teaching) is a popular summarization dataset that consists of news articles from CNN and DailyMail, paired with corresponding bullet-point summaries. Originally developed for question answering, it was later adapted for summarization (nallapati2016abstractive). For this dataset, we use pre-generated outputs from stiennon2020learning, which include results from 11 different systems, including human references and an extractive baseline. Since our meta-evaluation datasets for summarization contain inputs from CNN/DailyMail, we ensure there is no overlap in the source texts between these datasets when sampling the inputs.
  • Newsroom (grusky2018newsroom) is a large-scale dataset of news articles and their summaries, collected from diverse sources, domains, authors and time range. In addition to newly generated LLM outputs, we use outputs of the systems evaluated in the original paper, which include three summarization systems and two extractive baselines.
  • SAMSum dataset (gliwa2019samsum) addresses the dialogue summarization task, containing dialogues with short summaries created by linguists. In addition to outputs from several different LLMs, we use pre-generated outputs of six systems from (gao2022dialsummeval).
  • TL;DR (volske2017tl) is a collection of Reddit posts with user-created summaries. Unlike many popular datasets that focus on news articles, TL;DR contains informal and less structured texts spanning diverse topics. Similarly to CNN/DailyMail, we use outputs from stiennon2020learning.
  • XSum (narayan2018don) is a dataset for the extreme summarization task, designed for abstractive approaches. Unlike datasets such as Newsroom or CNN/DailyMail, which favor extractive summarization, XSum contains BBC articles paired with concise, single-sentence summaries. We utilize pre-generated outputs from five systems and baselines evaluated in the original paper, along with newly generated output from LLMs.
@@ A.1.2 Aspects
  • Accuracy is largely synonymous with consistency. It evaluates whether the factual information from the source is accurately represented in the summary. Following stiennon2020learning, we define accuracy as: Extent to which the factual information in the summary accurately matches the source text. An accurate summary should not contain information that is not present in the source text, should not contradict the source text, and generally should not be misleading.
@@
  • Coverage evaluates how much of the important information from the source text is covered by the summary. In this sense, it is closely related to relevance, however, it does not include non-redundancy as a criterion. We use a definition adapted from stiennon2020learning: Extent to which the summary covers the important information in the source text. A summary has good coverage if it mentions the main information from the source text that is important to understand the events described in the text. A summary has poor coverage if someone reading only the summary would be missing several important pieces of information about the event in the source text.
  • Coherence refers to the structural quality of the summary and involves attributes such as cohesion, consistency and relevance (reinhart1980conditions). It is determined by both semantic and formal structure of the text. We define coherence as: Extent to which the summary is well-structured and organized, presenting information in a logical order that flows naturally from sentence to sentence. Coherent summary forms a unified body of information and makes it easy to understand the main ideas.
@@
Table 5: Input data formats used in our dataset for data-to-text tasks.
@@ A.2 Data-to-text
For data-to-text category, we collected inputs from four datasets that represent four distinct tasks: table-to-text, RDF-to-text, attribute-value list to text and logical NLG. Since it is important for an NLG evaluation method to reliably evaluate outputs with respect to structured data in different formats, we include a number of different input formats in our training data, including JSON, CSV and linearized tables with markup. Table 5 provides an overview of the input formats. The list of evaluated systems in our dataset is shown in Table 7.
@@ A.2.1 Datasets
  • E2E NLG dataset (duvsek2020evaluating; novikova2017e2e) was chosen as a representative of simple data-to-text tasks. In this dataset, models are tasked with generating descriptions of restaurant venues based on attribute-value-based meaning representations (MRs). Target descriptions were crowdsourced using textual and pictorial representations of MRs as stimuli. To ensure diversity, we selected pre-generated outputs from five systems described by (duvsek2020evaluating), considering their model architectures and evaluation results across various metrics and aspects. These were extended by new outputs from several LLMs. The inputs and pre-generated outputs were sampled from the test set. Three different input formats are used in the training data (see Table 5).
  • WebNLG 2020 (ferreira20202020) is an RDF-to-text dataset designed for generating natural language text from RDF triples collected from DBPedia knowledge base (mendes2012dbpedia). Each input contains between one and seven triples, where each triple represents a binary relation in the form (subject, property, object). Similarly to E2E NLG, we selected a diverse set of output systems based on model architectures and evaluation results from the WebNLG+ 2020 Challenge, and used additional LLMs to generate new outputs.
  • ToTTo (parikh2020totto) is an open-domain table-to-text generation dataset. It consists of Wikipedia tables with highlighted cells, and the task is to generate single-sentence descriptions of the data in these highlighted cells. Inputs are provided in two formats: full tables, which include indices to highlighted cells, and linearized tables, where only the highlighted data is presented in a linear order, while the structure is annotated with markup tags. As we observed that even medium-sized open-weight LLMs often struggle with hallucinations when using full tables, we restricted our evaluation to linearized tables.
  • LogicNLG (chen2020logical) introduced the task of logical NLG, where models generate statements that can be logically entailed from the data in a table. This task involves various aggregations and comparisons, making it more difficult than simply transforming structured data to free-form text. Although the task explicitly requires generating five logical statements per input, we sampled between one to five statements from each generated output to increase the diversity of output lengths. Inputs were formatted as CSV with “|” as a separator. In addition to new LLM outputs, we used existing system outputs from zhao-etal-2023-investigating.
@@ A.2.2 Aspects
  • Faithfulness measures whether all information in the generated text is supported by the data, making it equivalent to precision. Similarly to factual consistency in summarization, we consider both factual and non-factual hallucinations as errors. For our evaluations, we define faithfulness as: Extent to which the information in the text is supported by the data.
@@
  • Fluency refers to the formal quality of the generated text, and includes grammaticality, naturalness and readability. Some definitions also include coherence (ferreira20202020), although coherence is usually treated as a separate aspect. For our evaluation, we define fluency as: Extent to which the text is grammatical, natural and easy to understand.
@@ A.3 Dialogue Response Generation
For dialogue response generation, we source the inputs from three dialogue datasets, focusing on open-domain non-task-oriented dialogue response generation. Table 8 provides an overview of all evaluated systems in the training dataset.
@@ A.3.1 Datasets
  • Wizard of Wikipedia (dinan2018wizard) consists of conversations grounded in one of 1365 topics and corresponding knowledge retrieved from Wikipedia. In these conversations, either participant may select the topic and initiate the discussion, although they have asymmetric roles. One participant takes on the role of the wizard, an expert with access to a topic-relevant knowledge, on which they can base their responses. The other participant acts as an apprentice, a curious learner that is eager to discuss the chosen topic. To create inputs of varying length, we randomly select a dialogue history length between two turns and the full conversation, and truncate the dialogue to this length. The last utterance is replaced by a system output, except when the reference is used as evaluated output. In addition to newly generated LLM responses, we include human responses from the original dataset in the outputs.
  • EmpatheticDialogues (rashkin2018towards) is a dataset of dialogues grounded in emotional situations, designed to train and evaluate dialogue models on empathetic response generation. Each conversation is associated with an emotional label, where one of the participants describes a situation in which they experienced a given emotion. We sample up to five turns from each dialogue and replace the last utterance with a generated response. The emotion label is used as additional context for annotator LLMs to evaluate the appropriateness and empathy of the responses, but is excluded from the prompts used for system output generation.
  • DailyDialog (li2017dailydialog) includes conversations on various daily life topics, annotated with emotion labels and communicative intents. We include data from DailyDialog to represent diverse topics and scenarios in the training set. Alongside newly generated responses, we also collect pre-generated outputs from three sources (gupta2019investigating; huang2020grade; zhao2020designing) to represent older dialogue systems.
@@ A.3.2 Aspects
  • Interestingness is concerned with the informational value of the response, specifically whether it presents stimulating ideas, facts or opinions. As it is one of the more subjective evaluation aspects, we define it vaguely and let the annotator models determine the criteria for interestingness: Extent to which the response is interesting given the dialogue history.
  • Engagingness is closely related to interestingness but is sometimes treated as a distinct aspect (e.g., mehri2020unsupervised; see-etal-2019-makes). While interestingness focuses on the context itself, engagingness emphasizes maintaining the user’s attention and encouraging them to continue with the conversation. For our purposes, engagingness is defined as: Extent to which the response captures and maintains the user’s interest, encouraging further interaction. Engaging responses contain opinions, preferences, thoughts or interesting facts.
@@ A.4 Story Generation
The NLG tasks discussed so far generally contain inputs that are relatively longer compared to the outputs. This pattern is especially common in tasks like summarization and dialogue generation, although certain data-to-text tasks also share this characteristic. To represent scenarios with short inputs and long outputs, we include story generation in the training data. Table 9 lists all evaluated systems in our dataset.
@@ A.4.1 Datasets
As the source of inputs, we use WritingPrompts (fan2018hierarchical), a story generation dataset derived from Reddit’s WritingPrompts subreddit, where users submit prompts that can inspire other users to write stories. The dataset consists of a diverse range of topics, story lengths and writing styles. We reuse existing outputs from the OpenMEVA dataset (guan2021openmeva) and generate additional outputs by four LLMs. To increase the diversity of generated stories in terms of their length, we generate the outputs with two different prompt versions, each requiring a different length of the story. The outputs are then randomly sampled from either the shorter or the longer set. As we observed a tendency of LLMs to generate a long list of errors for longer inputs, our evaluator models are instructed to limit the number of identified errors to a maximum of eight, and to prioritize the most severe ones if necessary.
@@ A.4.2 Aspects
While relevance and coherence are two commonly used aspects for story generation, there is no consensus on which other evaluation aspects are the most relevant. Inspired by social sciences, chhun2022human propose four additional aspects, aimed at providing a complete and non-redundant set of criteria. Following their work, we adopt the aspects defined in the HANNA benchmark, using our own definitions for most of them:
  • Relevance measures the degree to which a story aligns with the given prompt (chhun2022human; chiang2023can), title (jhamtani2020narrative; yao2019plan; xie2023next) or story beginning wang2020narrative. In some cases, relevance also evaluates whether the story remains on-topic for its duration (e.g., goldfarb2020content). Since our inputs are prompts, we define relevance as: Extent to which the story is relevant to the writing prompt.
  • Coherence in story generation typically refers to logical consistency and narrative flow (e.g., yao2019plan; li2023collaborative; jhamtani2020narrative). Other works define coherence more vaguely, such as how much the story “makes sense” (chhun2022human) or how well its sentences “fit together” (xie2023next). For our purposes, coherence is defined as: Extent to which the story is logically consistent and coherent.
  • Engagement is a subjective and often vaguely defined aspect that evaluates how engaging the story is to the reader (e.g., chhun2022human; li2023collaborative). Due to its inherent subjectivity, we apply a simple definition and leave its interpretation to the evaluator models: Extent to which the story is engaging and interesting.
@@ A.5 Question Answering
The question answering subset of our dataset consists of two distinct tasks: narrative question answering and table question answering. The inputs consist of a question and structured data in which the answer should be grounded. The evaluated systems in our training data are listed in Table 10.